# Low-Rank and Sparse Model Merging for Multi-Lingual Speech Recognition and Translation

### v2

@@ 5.2 Multi-Task Model Merging
| System | Avg. WER | Avg. BLEU |
| Pretrained | 19.88 | 25.48 |
| Finetuned | 19.05 | 26.18 |
| Multi-task training | 19.00 | 25.90 |
| Weight Averaging | 18.84 | 26.18 |
| Task Arithmetic | 18.76 | 26.30 |
| MP-Merging | 18.62 | 26.40 |
| SVP-Merging | 18.72 | 26.38 |
| LoRS-Merging | 18.45 | 26.48 |
System WER BLEU ca de es fr it Avg. ca de es fr it Avg. Pretrained 20.6 19.6 14.7 24.5 19.4 19.88 21.1 24.1 28.6 26.8 26.8 25.48 Finetuned 19.5 19.7 14.4 22.1 19.2 19.05 22.6 24.6 29.2 27.2 27.3 26.18 ML and MT training 20.5 19.7 14.6 24.5 19.4 19.86 21.3 24.3 28.3 27.1 26.9 25.58 ML and MT Task Arithmetic 18.9 19.2 14.1 23.7 18.4 18.96 22.2 24.4 29.0 27.3 26.9 25.96 ML and MT LoRS-Merging 18.7 19.1 14.0 23.8 18.0 18.82 22.2 24.8 29.0 27.5 27.0 26.10 MT training 17.0 19.7 14.4 24.2 19.4 19.00 22.3 24.6 28.7 27.0 26.9 25.90 + ML Task Arithmetic 18.1 19.0 14.2 24.5 20.6 19.37 22.7 24.7 28.6 27.3 26.5 25.96 + ML LoRS-Merging 18.1 19.0 14.1 24.2 20.3 19.23 22.4 24.5 29.1 27.6 26.7 26.06 ML training 17.1 21.8 15.1 22.6 21.9 19.69 21.4 24.4 28.8 26.8 27.2 25.72 + MT Task Arithmetic 17.1 18.5 13.3 22.7 18.0 18.00 22.6 25.0 29.2 27.5 26.9 26.24 + MT LoRS-Merging 16.9 18.3 13.3 22.4 17.8 17.82 22.8 25.2 29.3 27.6 27.0 26.38
@@
ASR and ST tasks for the same language can mutually benefit from each other: Table 3 presents the multi-task results with the high-resource language set. In general, multi-task training performs similarly to finetuned models on ASR but is a lot worse on ST. This is likely due to the substantial differences in optimal hyper-parameter configurations between the two tasks. Model merging methods clearly outperform finetuned models, which not only demonstrates their effectiveness but also shows the mutual benefits between ASR and ST. In terms of performance gains, the improvement in ASR is greater than in ST. We attribute this to the fact that ASR is inherently simpler than ST and can be viewed as a step in the ST task. Furthermore, as before, model merging methods combined with pruning further improve performance, and the proposed LoRS-Merging achieves the best performance across the table.
@@ 5.3 Multi-Lingual Multi-Task Model Merging
ML and MT training: Fine-tuning on all languages and both tasks jointly.
ML and MT merging: Fine-tuning on each language for each task separately and merging all.
MT training and ML merging: Fine-tuning both tasks jointly for each language, and merging models from different languages.
ML training and MT merging: Fine-tuning on all languages jointly for each task, and merging models from different tasks.
Table 4 displays the multi-lingual and multi-task results with the high-resource language set. Multi-lingual and multi-task training shows little improvement over the pretrained model, due to negative interference during training and the use of a unified training configuration for all languages and tasks. Nevertheless, the performance of multi-lingual and multi-task merging is on par with that of finetuned models, further underscoring the superiority of model merging. ML training followed by MT merging achieves the best performance, even significantly outperforming finetuned models. Although we did not observe the same phenomenon on the low-resource language set, this suggests the potential of using a combination of training and merging to achieve better performance. We provide additional experiments on the low-resource language set in Appendix B to demonstrate the robustness and generalisability of model merging and LoRS-Merging.
@@ 5.4 Effect of Numbers of Languages
LoRS-Merging improvements are consistent across different numbers of languages: As the number of languages increases, the performance of both TA and LoRS-Merging degrades due to negative interference between languages. LoRS-Merging consistently outperforms TA in both ASR and ST tasks, and even surpasses the finetuned models in the ASR task. This is likely due to LoRS-Merging further reducing model redundancy, therefore alleviating negative interference. Additionally, we observe that the optimal learning rate for the finetuned ASR model is significantly larger compared to the ST task. This may lead to overfitting, whereas LoRS-Merging improves model generalisation through model merging while reducing language interference, thus outperforming the finetuned models for the ASR task.
@@ 5.5 Effect of Language Data Scale
We then demonstrate the robustness of merging methods to different training data sizes for both tasks. Fig. 4 shows the WER (top) and BLEU (bottom) scores for ASR and ST at different data scales, respectively. As the data scale increases, the performance of multi-lingual training does not always improve. This may be because the pretrained model already performs well, and the significant language interference and conflict in multi-lingual training hinder the effective improvement of multi-language performance. Furthermore, the performance loss of model merging increases with data scale, compared to finetuned models. It can be explained by the fact that larger training data tends to increase the divergence in the optimisation trajectories of different finetuned models, resulting in the breakdown of linear mode connectivity, which leads to a greater performance loss. Moreover, LoRS-Merging still achieves obvious and stable improvement compared to TA.
@@ 6 Conclusion
This paper explores model merging for multi-lingual ASR and ST on pre-trained speech models and proposes the LoRS-Merging approach. LoRS-Merging combines low-rank and sparse pruning to retain essential structures, eliminate redundant parameters and mitigate language and task interference. Experiments across 10 languages show that LoRS-Merging effectively alleviates language interference and significantly outperforms multi-lingual multi-task training baselines.
@@ 7 Limitations
Appendix A Hyper-Parameter Details
The detailed hyper-parameter settings for each language are shown in Table 5 for ASR and Table 6 for ST, respectively.
@@ Appendix B Results of Low-Resource Language Set
The results of the low-resource language set are shown in this section. Specifically, Table 7 and 8 show the multi-lingual single task training and merging for ASR and ST respectively.
@@
Then, Table 9 shows the uni-lingual multi-task training and merging performance (c.f. compare to 3 for high-resource languages).
@@
Appendix C Detailed Results on Multi-task merging
System WER BLEU ca de es fr it Avg. ca de es fr it Avg. Pretrained 20.6 19.6 14.7 24.5 19.4 19.88 21.1 24.1 28.6 26.8 26.8 25.48 Finetuned 19.5 19.7 14.4 22.1 19.2 19.05 22.6 24.6 29.2 27.2 27.3 26.18 Multi-task training 17.0 19.7 14.4 24.2 19.4 19.00 22.3 24.6 28.7 27.0 26.9 25.90 Weight Averaging 17.1 19.6 13.9 23.7 19.6 18.84 22.9 24.4 29.0 27.7 26.9 26.18 Task Arithmetic 17.2 19.3 14.0 23.3 19.7 18.76 23.4 24.5 28.9 27.7 27.0 26.30 MP-Merging 17.8 19.5 14.4 23.8 17.2 18.62 23.1 24.5 29.1 27.9 27.4 26.40 SVP-Merging 18.0 19.4 14.4 23.7 17.7 18.72 22.9 24.7 29.1 27.8 27.4 26.38 LoRS-Merging 17.5 19.4 14.2 23.1 17.7 18.45 23.1 24.5 29.3 27.9 27.6 26.48

### v3

@@ 5.2 Multi-Task Model Merging
System Avg. WER Avg. BLEU Pretrained 19.88 - 25.48 - Finetuned 19.05 100.0% 26.18 100.0% Multi-task training 19.00 106.0% 25.90 60.0% Sequential training 18.95 112.0% 26.12 91.4% Weight Averaging 18.84 125.3% 26.18 100.0% Task Arithmetic 18.76 134.9% 26.30 117.1% TIES-Merging 18.60 154.2% 26.38 128.6% DARE 18.71 141.0% 26.28 114.3% TSV-M 18.70 142.2% 26.40 131.4% LoRS-Merging 18.39 179.5% 26.56 154.3%
System WER BLEU ca de es fr it Avg. ca de es fr it Avg. Pretrained 20.6 19.6 14.7 24.5 19.4 19.88 - 21.1 24.1 28.6 26.8 26.8 25.48 - Finetuned 19.5 19.7 14.4 22.1 19.2 19.05 100.0% 22.6 24.6 29.2 27.2 27.3 26.18 100.0% ML and MT training 20.5 19.7 14.6 24.5 19.4 19.86 2.4% 21.3 24.3 28.3 27.1 26.9 25.58 14.3% ML and MT Task Arithmetic 18.9 19.2 14.1 23.7 18.4 18.96 110.8% 22.2 24.4 29.0 27.3 26.9 25.96 68.6% ML and MT LoRS-Merging 18.7 19.1 14.0 23.8 18.0 18.82 127.7% 22.2 24.8 29.0 27.5 27.0 26.10 88.6% MT training 17.0 19.7 14.4 24.2 19.4 19.00 - 22.3 24.6 28.7 27.0 26.9 25.90 - + ML Task Arithmetic 18.1 19.0 14.2 24.5 20.6 19.37 61.4% 22.7 24.7 28.6 27.3 26.5 25.96 68.6% + ML LoRS-Merging 18.1 19.0 14.1 24.2 20.3 19.23 78.3% 22.4 24.5 29.1 27.6 26.7 26.06 82.9% ML training 17.1 21.8 15.1 22.6 21.9 19.69 - 21.4 24.4 28.8 26.8 27.2 25.72 - + MT Task Arithmetic 17.1 18.5 13.3 22.7 18.0 18.00 226.5% 22.6 25.0 29.2 27.5 26.9 26.24 108.6% + MT LoRS-Merging 16.9 18.3 13.3 22.4 17.8 17.82 248.2% 22.8 25.2 29.3 27.6 27.0 26.38 128.6%
@@
ASR and ST tasks for the same language can mutually benefit from each other: Table 3 presents the multi-task results with the high-resource language set. In general, multi-task training performs similarly to finetuned models on ASR but is a lot worse on ST. This is likely due to the substantial differences in optimal hyper-parameter configurations between the two tasks. Sequential training performs similarly to finetuned models overall, as it also benefits from training independence. Model merging methods clearly outperform finetuned models, which not only demonstrates their effectiveness but also shows the mutual benefits between ASR and ST. In terms of performance gains, the improvement in ASR is greater than in ST. We attribute this to the fact that ASR is inherently simpler than ST and can be viewed as a step in the ST task. Furthermore, as before, model merging methods combined with pruning further improve performance, and the proposed LoRS-Merging achieves the best performance across the table.
@@ 5.3 Multi-Lingual Multi-Task Model Merging
ML and MT training: Finetuning on all languages and both tasks jointly.
ML and MT merging: Finetuning on each language for each task separately and merging all.
MT training and ML merging: Finetuning both tasks jointly for each language, and merging models from different languages.
ML training and MT merging: Finetuning on all languages jointly for each task, and merging models from different tasks.
Table 4 displays the multi-lingual and multi-task results with the high-resource language set. Multi-lingual and multi-task training shows little improvement over the pretrained model, due to language interference during training and the use of a unified training configuration for all languages and tasks. Nevertheless, the performance of multi-lingual and multi-task merging is on par with that of finetuned models, further underscoring the superiority of model merging. ML training followed by MT merging achieves the best performance, even significantly outperforming finetuned models. Although we did not observe the same phenomenon on the low-resource language set, this suggests the potential of using a combination of training and merging to achieve better performance. We provide additional experiments on the low-resource language set in Appendix B to demonstrate the robustness and generalisability of model merging and LoRS-Merging.
@@ 5.4 Effect of Numbers of Languages
LoRS-Merging improvements are consistent across different numbers of languages: As the number of languages increases, the performance of both TA and LoRS-Merging degrades due to negative interference between languages. LoRS-Merging consistently outperforms TA in both ASR and ST tasks. Notably, in the ASR task, it even surpasses the finetuned models. This is primarily because finetuned models contain substantial redundancy (see Fig. 5), whereas LoRS-Merging reduces redundancy through pruning, leading to significant performance improvements. Additionally, we observe that the optimal learning rate for the finetuned ASR model is significantly larger compared to the ST task. This may lead to overfitting in ASR. In contrast, LoRS-Merging improves generalisation through model merging, thus outperforming the finetuned models for the ASR task.
@@ 5.5 Effect of Language Data Scale
We then demonstrate the robustness of merging methods to different training data sizes for both tasks. Fig. 4 shows the WER (top) and BLEU (bottom) scores for ASR and ST at different data scales, respectively. As the data scale increases, the performance of multi-lingual training does not always improve. This is because the multi-lingual capabilities of pretrained models are already near convergence, and only meticulous training can further improve performance. Increased training data amplifies both language interference and the negative effects of uniform training configurations, thereby offsetting the gains from increased data. Furthermore, the performance loss of model merging increases with data scale, compared to finetuned models. It can be explained by the fact that larger training data tends to increase the divergence in the optimisation trajectories of different finetuned models, resulting in the breakdown of linear mode connectivity, which leads to a greater performance loss. Moreover, LoRS-Merging still achieves obvious and stable improvement compared to TA.
@@ 6 Conclusion
This paper explores model merging for multi-lingual ASR and ST on pretrained speech models and proposes the LoRS-Merging approach. LoRS-Merging combines low-rank and sparse pruning to retain essential structures and reduce redundant parameters. Experiments across 10 languages show that LoRS-Merging effectively alleviates language interference and significantly outperforms multi-lingual multi-task training, sequential training, and other merging methods.
@@ 7 Limitations
Appendix A Details of the Experimental Setup
For multi-lingual and multi-task training, a uniform training configuration is used across all languages and tasks. For sequential training, considering that there are 5! = 120 possible sequences for 5 languages, and the optimal training configuration for the same language differs across sequences, the hyper-parameter search cost for sequential training is much higher than that for the model merging. To simplify the configuration, we select 5 sequences for the experiments, corresponding to all cyclic permutations of the language order ca-de-es-fr-it, and report the results from the sequence that yields the best average performance.
MP and SVP are applied to each linear layer. The detailed hyper-parameter settings for each language are shown in Table 5 for ASR and Table 6 for ST, respectively.
@@ Appendix B Results of Low-Resource Language Set
The results of the low-resource language set are shown in this section. Specifically, Table 7 and 8 show the multi-lingual single-task training and merging for ASR and ST respectively.
@@
Then, Table 9 shows the multi-task single-language training and merging performance (c.f. compare to Table 3 for high-resource languages).
@@
Appendix C Detailed Results on Multi-Task Model Merging
System WER BLEU ca de es fr it Avg. ca de es fr it Avg. Pretrained 20.6 19.6 14.7 24.5 19.4 19.88 21.1 24.1 28.6 26.8 26.8 25.48 Finetuned 19.5 19.7 14.4 22.1 19.2 19.05 22.6 24.6 29.2 27.2 27.3 26.18 Multi-task training 17.0 19.7 14.4 24.2 19.4 19.00 22.3 24.6 28.7 27.0 26.9 25.90 Sequential training 16.7 19.4 14.3 24.6 19.4 18.95 22.9 24.8 28.7 27.3 26.9 26.12 Weight Averaging 17.1 19.6 13.9 23.7 19.6 18.84 22.9 24.4 29.0 27.7 26.9 26.18 Task Arithmetic 17.2 19.3 14.0 23.3 19.7 18.76 23.4 24.5 28.9 27.7 27.0 26.30 TIES-Merging 17.7 19.5 14.4 23.6 17.4 18.60 23.1 24.5 29.1 27.7 27.5 26.38 DARE 17.5 19.4 14.2 23.5 18.6 18.71 23.2 24.5 29.0 27.6 27.1 26.28 TSV-M 17.8 19.4 14.3 23.7 17.9 18.70 23.0 24.7 29.2 27.7 27.4 26.40 LoRS-Merging 17.3 19.4 14.1 23.1 17.7 18.39 23.3 24.6 29.3 28.0 27.6 26.56